# EVM-Fusion: An Explainable Vision Mamba Architecture with Neural Algorithmic Fusion

### v3

@@ 3.3 Core Fusion Module
This module intelligently integrates the feature vectors $v_{D},v_{U},v_{T}$.
@@ 3.3.1 Stage A: Cross-Modal Context Encoding
The input feature vectors are stacked to form $v_{in}=[v_{D};v_{U};v_{T}]\in\mathbb{R}^{N_{paths}\times d_{model\_fusion}}$, where $N_{paths}=3$. Then $V_{in}$ is processed by an MHA layer acting as a cross-modal attention mechanism.
$AttendedFeatures,W_{CMA}=MHA(v_{in})$ (9)
The attention weights $W_{CMA}\in\mathbb{R}^{N_{paths}\times N_{paths}}$ (averaged over heads) indicate the learned inter-path dependencies and are saved as Cross-Modal Attn Weights for XAI. Then a residual connection followed by Layer Normalization is applied.
$\begin{split}v_{contextual}=LayerNorm(v_{in}+\\ AttendedFeatures)\end{split}$ (10)
@@ 3.3.2 Stage B: Neural Algorithmic Fusion (NAF) Block
The NAF block takes the contextualized vector $v_{contextual}$ and learns an iterative algorithm to fuse them.
The initial input to the NAF controller $c_{0}=W_{init\_ctrl}v_{contextual}+b_{init\_ctrl}$, where $c_{0}\in\mathbb{R}^{d_{primitive\_out}}$
The controller, a Gated Recurrent Unit (GRU) [44], takes the previous fused state $S_{fused}^{(k-1)}$ and its own previous hidden state $h_{ctrl}^{(k-1)}$ to produce an output $o_{ctrl}^{(k)}$ and new hidden state $h_{ctrl}^{(k)}$
$z_{k}=\sigma(W_{z}S_{fused}^{(k-1)}+U_{z}h_{ctrl}^{(k-1)}+b_{z})$ (11)
$r_{k}=\sigma(W_{r}S_{fused}^{(k-1)}+U_{r}h_{ctrl}^{(k-1)}+b_{r})$ (12)
$\tilde{h}_{ctrl}^{(k)}=tanh(W_{h}S_{fused}^{(k-1)}+U_{h}(r_{k}\odot h_{ctrl}^{(k-1)})+b_{h})$ (13)
${h}_{ctrl}^{(k)}=(1-z_{k})\odot h_{ctrl}^{(k-1)}+z_{k}\odot\tilde{h}_{ctrl}^{(k)}$ (14)
actually $o_{ctrl}^{(k)}$ is typically ${h}_{ctrl}^{(k)}$ and the iterative fused state $S_{fused}^{(0)}$ is $c_{0}$. The iteration process will be performed $K_{NAF}$ times
The controller output $o_{ctrl}^{(k)}$ is used to generate mixing weights $\alpha^{(k)}\in\mathbb{R}^{N_{primitives}}$ for the primitive operations:
$\alpha^{(k)}=softmax(W_{mix}o_{ctrl}^{(k)}+b_{mix})$ (15)
A bank of $N_{primitives}$ distinct primitive operations $\mathcal{P}=\{Prim_{1},...,Prim_{N_{primitives}}\}$ (each typically a MLP in my implementation) processes the static $v_{contextual}$:
$P_{j}^{(k)}=Prim_{j}(v_{contextual})\in\mathbb{R}^{d_{primitive\_out}}$ (16)
and are mixed using $\alpha^{(k)}$:
$S_{mix}^{(k)}=\sum_{j=1}^{N_{primitives}}\alpha^{(k)}P_{j}^{(k)}$ (17)
During the iteration, I update the state using residual connection: If $k=1,S_{fused}^{(1)}=S_{mix}^{(1)}$. If $k>1,S_{fused}^{(k)}=LayerNorm(S_{fused}^{(k-1)}+S_{mix}^{(k)})$
The final output of the NAF block $v_{NAF}$, is obtained by further processing $S_{fused}^{(K_{NAF})}$ through a linear layer and Layer-norm after $K_{NAF}$ iterations.
The NAF block’s strength lies in its learned, iterative refinement. The controller (GRU) acts as a policy network, dynamically determining at each step k how to best combine the outputs of various general-purpose primitive functions ($\text{Prim}_{j}$) based on the current state of fusion ($S_{fused}^{(k-1)}$). This allows for a more adaptive and potentially deeper fusion process than a fixed set of operations. The primitives always operating on the initial comprehensive feature set ($v_{contextual}$) ensures access to rich information, while the iterative state $S_{fused}^{(k)}$ allows the controller to track the fusion progress. Residual connections and LayerNorm aid in stabilizing training for multiple iterations.
Table 2: Detailed Comparison of Model Performance Metrics with Per-Class Scores. Best values in each column are in red, second best in blue.
@@
EVM-DUHF 0.9975 0.9975 0.9975 0.9975 0.9858 1.0000 0.9928 1.0000 0.9922 0.9961 1.0000 0.9917 0.9959 1.0000 1.0000 1.0000 1.0000 0.9967 0.9983 1.0000 1.0000 1.0000 0.9957 0.9957 0.9957 1.0000 1.0000 1.0000 0.9933 1.0000 0.9966
EVM-DHF 0.8820 0.8907 0.8820 0.8809 0.7193 0.7885 0.7523 0.5926 0.5000 0.5424 0.8269 0.9669 0.8914 0.9958 0.9754 0.9855 0.9934 1.0000 0.9967 0.9761 0.9879 0.9820 0.9936 0.6667 0.7980 0.9887 0.9887 0.9887 0.6304 0.7838 0.6988
EVM-DU 0.8934 0.9028 0.8934 0.8945 0.7556 0.8173 0.7852 0.6607 0.5781 0.6167 0.8776 0.8884 0.8830 0.9877 0.9877 0.9877 0.9967 0.9933 0.9950 0.9879 0.9839 0.9859 0.9773 0.7350 0.8390 0.9887 0.9925 0.9906 0.6158 0.8446 0.7123
EVM-UHF 0.7685 0.8291 0.7685 0.7606 0.3701 0.7740 0.5008 0.4471 0.2969 0.3568 0.9069 0.9256 0.9162 0.9138 0.2172 0.3510 0.9700 0.9700 0.9700 0.9194 0.9194 0.9194 0.9706 0.7051 0.8168 0.9630 0.9811 0.9720 0.6373 0.8784 0.7386
DenseNet[25] 0.8969 0.8986 0.8969 0.8954 0.7021 0.7933 0.7449 0.5743 0.4531 0.5066 0.8613 0.9752 0.9147 0.9835 0.9754 0.9794 1.0000 0.9967 0.9983 0.9679 0.9718 0.9698 0.9948 0.8120 0.8941 0.9813 0.9887 0.9850 0.7547 0.8108 0.7818
U-Net[27] 0.8374 0.8471 0.8374 0.8378 0.6937 0.5337 0.6033 0.4489 0.6172 0.5197 0.8206 0.8884 0.8532 0.9424 0.9385 0.9405 0.9637 0.9733 0.9685 0.9469 0.8629 0.9030 0.8956 0.6966 0.7837 0.9059 0.9811 0.9420 0.7079 0.8514 0.7730
U-DenseNet[28] 0.8537 0.8526 0.8537 0.8495 0.6484 0.7981 0.7155 0.4750 0.2969 0.3654 0.8297 0.9463 0.8842 0.9732 0.8934 0.9316 0.9767 0.9767 0.9767 0.8716 0.9032 0.8871 0.9439 0.7906 0.8605 0.9348 0.9736 0.9538 0.7303 0.7500 0.7400
ViT[32] 0.8969 0.9018 0.8969 0.8958 0.7149 0.8317 0.7689 0.6316 0.4688 0.5381 0.8988 0.9545 0.9259 0.9917 0.9754 0.9835 0.9933 0.9900 0.9917 0.9683 0.9839 0.9760 0.9836 0.7692 0.8633 0.9849 0.9849 0.9849 0.6793 0.8446 0.7530
VMamba[42] 0.9033 0.9073 0.9033 0.9003 0.7160 0.8846 0.7914 0.6962 0.4297 0.5314 0.8745 0.9793 0.9240 0.9877 0.9877 0.9877 1.0000 0.9900 0.9950 0.9837 0.9758 0.9798 0.9836 0.7692 0.8633 0.9778 0.9962 0.9869 0.7176 0.8243 0.7673
ConvNeXt[45] 0.9023 0.9056 0.9023 0.9010 0.7125 0.8221 0.7634 0.6146 0.4609 0.5268 0.8988 0.9545 0.9259 0.9918 0.9918 0.9918 1.0000 0.9933 0.9967 0.9799 0.9839 0.9819 0.9840 0.7906 0.8768 0.9850 0.9887 0.9868 0.7151 0.8649 0.7829
@@ 4 Experiments
There are three datasets in total. One is the brain tumor MRI dataset from Kaggle [46], which contains four categories: healthy, glioma, meningioma, and pituitary. Another is the breast cancer X-ray dataset from Kaggle [47], which contains two categories: 0, 1, where 0 represents benign and 1 represents malignant. The last one is the chest X-ray dataset from Mendeley Data [48] proposed by Daniel S. Kermany et al. [49], which contains three categories: normal, bacteria, virus, where normal represents health, bacteria represents bacterial pneumonia, and virus represents viral pneumonia. The specific distribution is shown in Table 1.
All images are uniformly resized to a standard size of 224 × 224 pixels. A variety of data augmentation measures were taken for the images in the training set, including rotation, translation, scaling, flipping, and brightness adjustment.The models are trained on the training set and validated with early stopping on the test set. The confusion matrix is obtained on the test set, as shown in Fig. 2. Based on the confusion matrix, other evaluation indicators are calculated, as shown in Table 2.
@@ 5 Discussion
The EVM-Fusion model demonstrates strong classification performance, underscoring its potential for robust multi-organ medical image analysis. Beyond quantitative metrics, a central aim was achieving intrinsic interpretability. This section discusses insights from the model’s built-in eXplainable AI (XAI) mechanisms, analyzing its behavior on correctly classified (Fig. 4) and misclassified (Fig. 3) samples to elucidate its operational dynamics and the benefits of its architecture.
Figure 3: Explainable AI (XAI) Analysis of Representative Misclassified Test Samples by EVM-DUHF. This figure presents a consolidated view of intrinsic explainability outputs for five misclassified samples from the 9-class test set. Noted from (a) to (e). Each sample is analyzed using eight distinct XAI visualizations. From top to bottom, from left to right, they are DenseNet Vision Mamba Forward $\Delta$ Map, DenseNet Vision Mamba Backward $\Delta$ Map, U-Net Vision Mamba Forward $\Delta$ Map, U-Net Vision Mamba Backward $\Delta$ Map, DenseNet Path Spatial Attention, U-Net Path Spatial Attention, Traditional Feature Importance (SE Attention bar plot), Cross-Modal Attention Weights (heatmap). These visualizations aim to elucidate the model’s internal feature prioritization and fusion dynamics in instances of incorrect classification, providing insights into potential failure modes and areas of confusion.
Vision Mamba’s Adaptive Processing ($\Delta$-value Maps): The $\Delta$-value maps from the Vision Mamba modules reveal dynamic, input-dependent information gating. Higher $\Delta$ values (warmer colors) signify a greater emphasis on current patch information, indicating regions of novelty or high salience. Conversely, lower $\Delta$ values (cooler colors) suggest reliance on previously propagated contextual states, often where information is redundant or predictable. For instance, in a correctly classified ”normal” chest X-ray (Fig. 4g), the DenseNet path’s Vim $\Delta$ maps show adaptive focus: the forward scan emphasizes lung boundaries upon encountering them, while the backward scan might highlight different boundary segments. The U-Net path’s Vim modules on the same sample might show more detailed, localized $\Delta$ peaks within the lung fields. Comparing forward and backward $\Delta$ maps (e.g., Fig. 4b, a malignant tumor) often reveals differing ”novelty” assessments based on the directional context of Mamba’s selective scan, showcasing its capacity for nuanced contextual understanding.
Path-Specific Spatial Attention: The spatial attention maps for the DenseNet and U-Net Mamba paths highlight prioritized image regions before global feature aggregation. In a correctly identified malignant tumor (Fig. 4b), these maps focus on the breast tissue, with DenseNet potentially emphasizing broader areas while U-Net targets more specific regions like the nipple, possibly indicating sensitivity to microcalcifications or lesion-dense areas. Interestingly, Fig. 4b also reveals a potential U-Net flaw, with some attention on blank image edges, underscoring the value of multi-path fusion to mitigate individual path imperfections.
Traditional Feature Importance (SE-Attention): The Squeeze-and-Excitation (SE) attention weights within the traditional feature path quantify the contribution of specific GLCM and LBP features. For a correctly classified ”normal” lung (Fig. 4g), features indicative of textural uniformity (LBP_bin_5 and LBP_bin_6) are often highly weighted. This demonstrates the model’s ability to leverage well-understood handcrafted features that complement deep learned representations, capturing intrinsic textural properties relevant to the diagnostic class.
Cross-Modal Attention: A cornerstone of EVM-Fusion’s adaptability is the dynamic nature of its cross-modal attention weights, which vary significantly across different input images (Fig. 4). For the benign tumors (Fig. 4a), all paths show increased attention towards the U-Net path, suggesting a reliance on its detailed local information for tumor characterization. Conversely, for a ”healthy” brain MRI (Fig. 4e), the attention shift, with U-Net drawing more from the traditional path if specific textures are key. This input-dependent, flexible weighting allows EVM-Fusion to adaptively modulate the influence of each feature pathway prior to the final Neural Algorithmic Fusion (NAF) stage. Notably, across many samples in Fig. 4, most paths often receive substantial attention (e.g., ¿25-30%), indicating their collective importance and complementary roles. This inherent dynamism and balanced contribution underpin the superior performance of the full 3-path EVM-Fusion (EVM-DUHF) over ablated versions lacking one or more paths, as it allows the model to maximize information gain from diverse feature types and reduce over-reliance on any single potentially fallible source.
@@
Figure 4: Explainable AI (XAI) Analysis of Representative Correctly Classified Test Samples by EVM-DUHF Across All Classes. This figure showcases the intrinsic explainability outputs for nine correctly classified samples, one from each of the distinct medical image classes in the test set. Noted from (a) to (i). Similar to Fig. 3, each sample is depicted through eight XAI visualizations. The analysis of these diverse, correctly classified instances aims to demonstrate the model’s adaptive behavior, the varying contributions of its different pathways, and its ability to focus on salient, class-relevant features across multiple organ types and imaging modalities.
@@
For (a) in Fig.3, both DenseNet and U-Net only focus on the same area of the breast. The model judges the malignant tumor (1) as benign (0), which indicates that there may be no tumor in this area. For (b) and (d) in Fig.3, it can be found that the DenseNet path focuses on the lower right corner, and U-Net pays almost no attention to the human body. This causes the model to even make cross-category errors, mistaking photos of the brain and chest for breasts. And (c) and (e) in Fig.3 also make the same mistake. Neither DenseNet nor U-Net pays good attention to the key organ - the lungs.

### v4

@@ 3.3 Core Fusion Module
This module intelligently integrates the feature vectors $\mathbf{v}_{D},\mathbf{v}_{U},\mathbf{v}_{T}$.
@@ 3.3.1 Stage A: Cross-Modal Context Encoding
To enable the different feature streams to inform one another before fusion, I first stack them into a matrix $\mathbf{v}_{in}=[\mathbf{v}_{D};\mathbf{v}_{U};\mathbf{v}_{T}]\in\mathbb{R}^{N_{paths}\times d_{model\_fusion}}$, where $N_{paths}=3$. Then $V_{in}$ is processed by an MHA layer acting as a cross-modal attention mechanism.
$\mathbf{V}_{\text{attended}},\mathbf{W}_{\text{CMA}}=\text{MHA}(\mathbf{V}_{\text{in}},\mathbf{V}_{\text{in}},\mathbf{V}_{\text{in}})$ (5)
The attention weights $\mathbf{W}_{\text{CMA}}\in\mathbb{R}^{N_{paths}\times N_{paths}}$ are saved for XAI. A residual connection followed by Layer Normalization is applied to produce the contextualized feature matrix:
$\mathbf{V}_{\text{contextual}}=\text{LayerNorm}(\mathbf{V}_{\text{in}}+\mathbf{V}_{\text{attended}})$ (6)
@@ 3.3.2 Stage B: Neural Algorithmic Fusion (NAF) Block
To move beyond a single-step fusion, the NAF block is designed to learn an iterative refinement algorithm. Its strength lies in a controller-driven process that dynamically mixes the outputs of several simple ”primitive” operations.
The process is initialized for $K_{\text{NAF}}$ iterations. At each step $k$, a controller (a GRU [44]) receives the current fused state $\mathbf{S}_{fused}^{(k-1)}$ and generates mixing weights $\alpha^{(k)}$ for a bank of $N_{primitives}$ primitive operations $\{\mathcal{P}_{j}\}$. Each primitive (a simple MLP) processes the static, comprehensive feature set $\mathbf{V}_{\text{contextual}}$. The outputs of the primitives are then mixed according to $\alpha^{(k)}$ to produce an update, which is added to the current state via a residual connection.
Algorithm 1 Neural Algorithmic Fusion (NAF) Block
1:A set of contextualized feature vectors $\{\mathbf{v}_{i}\}_{i=1}^{N_{\text{paths}}}$, where $\mathbf{v}_{i}\in\mathbb{R}^{d_{\text{model\_fusion}}}$.
2:The final fused feature vector $\mathbf{v}_{\text{NAF}}\in\mathbb{R}^{d_{\text{model\_fusion}}}$.
[…]
6:for $j=1\to N_{\text{primitives}}$ do $\triangleright$ Pre-compute primitives from static input
7:  $\mathbf{p}_{j}\leftarrow\text{Primitive}_{j}(\mathbf{v}_{\text{cat}})$ $\triangleright$ Each $\text{Primitive}_{j}$ is a small MLP; $\mathbf{p}_{j}\in\mathbb{R}^{d_{\text{state}}}$
8:end for
9:for $k=1\to K_{\text{NAF}}$ do $\triangleright$ Iteratively refine the fused state
10:  $\mathbf{h}_{\text{ctrl}}\leftarrow\text{GRU}(\mathbf{s}_{\text{fused}},\mathbf{h}_{\text{ctrl}})$ $\triangleright$ Controller observes current state and updates
11:  $\mathbf{\alpha}\leftarrow\text{Softmax}(\text{Linear}(\mathbf{h}_{\text{ctrl}}))$ $\triangleright$ Generate mixing weights $\mathbf{\alpha}\in\mathbb{R}^{N_{\text{primitives}}}$
12:  $\mathbf{s}_{\text{mix}}\leftarrow\sum_{j=1}^{N_{\text{primitives}}}\alpha_{j}\mathbf{p}_{j}$ $\triangleright$ Calculate weighted sum of primitives; $\mathbf{s}_{\text{mix}}\in\mathbb{R}^{d_{\text{state}}}$
13:  $\mathbf{s}_{\text{fused}}\leftarrow\text{LayerNorm}(\mathbf{s}_{\text{fused}}+\mathbf{s}_{\text{mix}})$ $\triangleright$ Update state with residual connection
14:end for
[…]
Table 2: Ablation results: Detailed comparison of performance metrics and per-category scores after removing some modules from the model. Best values in each column are in red, second best in blue.
@@
(Proposed Method) EVM-Fusion (DUHF) 0.9479 0.9557 0.9479 0.9484 0.9351 0.6923 0.7956 0.6484 0.9219 0.7613 0.9747 0.9545 0.9645 1.0000 1.0000 1.0000 0.9967 1.0000 0.9983 1.0000 0.9960 0.9980 0.9576 0.9658 0.9617 1.0000 1.0000 1.0000 0.9073 0.9257 0.9164
(Path Ablation) EVM-Fusion (DHF) 0.8820 0.8907 0.8820 0.8809 0.7193 0.7885 0.7523 0.5926 0.5000 0.5424 0.8269 0.9669 0.8914 0.9958 0.9754 0.9855 0.9934 1.0000 0.9967 0.9761 0.9879 0.9820 0.9936 0.6667 0.7980 0.9887 0.9887 0.9887 0.6304 0.7838 0.6988
(Path Ablation) EVM-Fusion (DU) 0.8934 0.9028 0.8934 0.8945 0.7556 0.8173 0.7852 0.6607 0.5781 0.6167 0.8776 0.8884 0.8830 0.9877 0.9877 0.9877 0.9967 0.9933 0.9950 0.9879 0.9839 0.9859 0.9773 0.7350 0.8390 0.9887 0.9925 0.9906 0.6158 0.8446 0.7123
(Path Ablation) EVM-Fusion (UHF) 0.7685 0.8291 0.7685 0.7606 0.3701 0.7740 0.5008 0.4471 0.2969 0.3568 0.9069 0.9256 0.9162 0.9138 0.2172 0.3510 0.9700 0.9700 0.9700 0.9194 0.9194 0.9194 0.9706 0.7051 0.8168 0.9630 0.9811 0.9720 0.6373 0.8784 0.7386
(Path Ablation) DenseNet[29] 0.8969 0.8986 0.8969 0.8954 0.7021 0.7933 0.7449 0.5743 0.4531 0.5066 0.8613 0.9752 0.9147 0.9835 0.9754 0.9794 1.0000 0.9967 0.9983 0.9679 0.9718 0.9698 0.9948 0.8120 0.8941 0.9813 0.9887 0.9850 0.7547 0.8108 0.7818
(Path Ablation) U-Net[18] 0.8374 0.8471 0.8374 0.8378 0.6937 0.5337 0.6033 0.4489 0.6172 0.5197 0.8206 0.8884 0.8532 0.9424 0.9385 0.9405 0.9637 0.9733 0.9685 0.9469 0.8629 0.9030 0.8956 0.6966 0.7837 0.9059 0.9811 0.9420 0.7079 0.8514 0.7730
(Path Ablation) U-DenseNet[19] 0.8537 0.8526 0.8537 0.8495 0.6484 0.7981 0.7155 0.4750 0.2969 0.3654 0.8297 0.9463 0.8842 0.9732 0.8934 0.9316 0.9767 0.9767 0.9767 0.8716 0.9032 0.8871 0.9439 0.7906 0.8605 0.9348 0.9736 0.9538 0.7303 0.7500 0.7400
(Fusion Ablation) Simple-Mean 0.8880 0.8930 0.8880 0.8869 0.6933 0.7933 0.7399 0.5657 0.4375 0.4934 0.8902 0.9380 0.9135 0.9758 0.9918 0.9837 0.9966 0.9900 0.9933 0.9757 0.9718 0.9737 0.9721 0.7436 0.8426 0.9924 0.9849 0.9886 0.6737 0.8649 0.7574
(Fusion Ablation) Simple-Concat 0.8721 0.8662 0.8721 0.8531 0.6211 0.9615 0.7547 0.4667 0.0547 0.0979 0.8014 0.9669 0.8764 0.9785 0.9344 0.9560 0.9900 0.9867 0.9883 0.9469 0.9355 0.9412 0.9412 0.8205 0.8767 0.9429 0.9962 0.9688 0.8346 0.7162 0.7709
(Fusion Ablation) NAF-Only 0.8681 0.8799 0.8681 0.8682 0.5860 0.8029 0.6775 0.5233 0.3516 0.4206 0.8731 0.9380 0.9044 0.9916 0.9713 0.9814 1.0000 0.8900 0.9418 0.9683 0.9839 0.9760 0.9778 0.7521 0.8502 0.9887 0.9925 0.9906 0.6868 0.8446 0.7576
(Fusion Ablation) CMA-Only 0.8934 0.9041 0.8934 0.8954 0.7500 0.6346 0.6875 0.5276 0.6719 0.5911 0.9036 0.9298 0.9165 0.9879 1.0000 0.9939 0.9966 0.9900 0.9933 0.9918 0.9798 0.9858 0.9892 0.7821 0.8735 0.9887 0.9887 0.9887 0.6878 0.8784 0.7715
The controller’s state update at step k follows standard GRU equations:
$\mathbf{h}_{\text{ctrl}}^{(k)}=\text{GRU}(\mathbf{S}_{\text{fused}}^{(k-1)},\mathbf{h}_{\text{ctrl}}^{(k-1)})$ (7)
[…]
$\mathbf{\alpha}^{(k)}=\text{softmax}(\mathbf{W}_{\text{mix}}\mathbf{h}_{\text{ctrl}}^{(k)}+\mathbf{b}_{\text{mix}})$ (8)
A bank of primitives processes $\mathbf{V}_{\text{contextual}}$ (here flattened and concatenated), and their outputs are mixed:
$\mathbf{S}_{\text{mix}}^{(k)}=\sum_{j=1}^{N_{\text{primitives}}}\alpha_{j}^{(k)}\mathcal{P}_{j}(\mathbf{V}_{\text{contextual}})$ (9)
The fused state is then updated with a residual connection:
$\mathbf{S}_{\text{fused}}^{(k)}=\text{LayerNorm}(\mathbf{S}_{\text{fused}}^{(k-1)}+\mathbf{S}_{\text{mix}}^{(k)})$ (10)
After $K_{NAF}$ iterations, the final state $\mathbf{S}_{\text{fused}}^{(K_{NAF})}$ is processed by a linear layer to produce the final fused vector $\mathbf{V}_{\text{NAF}}$.
Table 3: Detailed Comparison of Model Performance Metrics with Per-Class Scores and Parameters. Best values are in red, second best in blue.
Model Params (M) Overall 0 1 Bacteria Glioma Healthy Meningioma Normal Pituitary Virus
Acc P R F1 P R F1 P R F1 P R F1 P R F1 P R F1 P R F1 P R F1 P R F1 P R F1
(Proposed Method) EVM-Fusion 134.28 0.9479 0.9557 0.9479 0.9484 0.9351 0.6923 0.7956 0.6484 0.9219 0.7613 0.9747 0.9545 0.9645 1.0000 1.0000 1.0000 0.9967 1.0000 0.9983 1.0000 0.9960 0.9980 0.9576 0.9658 0.9617 1.0000 1.0000 1.0000 0.9073 0.9257 0.9164
Swim Transformer[31] 195.01 0.8136 0.8215 0.8136 0.8146 0.6859 0.6298 0.6566 0.4690 0.5312 0.4982 0.8359 0.9050 0.8690 0.8996 0.8811 0.8903 0.9801 0.9867 0.9834 0.8142 0.8306 0.8224 0.8864 0.6667 0.7610 0.9046 0.8943 0.8994 0.6043 0.7635 0.6746
Internimage[45] 221.43 0.8934 0.9028 0.8934 0.8945 0.7556 0.8173 0.7852 0.6607 0.5781 0.6167 0.8776 0.8884 0.8830 0.9877 0.9877 0.9877 0.9967 0.9933 0.9950 0.9879 0.9839 0.9859 0.9773 0.7350 0.8390 0.9887 0.9925 0.9906 0.6158 0.8446 0.7123
DeiT III[46] 303.88 0.9068 0.9105 0.9068 0.9054 0.7593 0.8798 0.8151 0.7368 0.5469 0.6278 0.8298 0.9669 0.8931 0.9959 0.9918 0.9938 1.0000 0.9867 0.9933 0.9880 0.9919 0.9899 0.9840 0.7906 0.8768 0.9888 1.0000 0.9944 0.6968 0.7297 0.7129
MambaVision[47] 226.41 0.9033 0.9073 0.9033 0.9003 0.7160 0.8846 0.7914 0.6962 0.4297 0.5314 0.8745 0.9793 0.9240 0.9877 0.9877 0.9877 1.0000 0.9900 0.9950 0.9837 0.9758 0.9798 0.9836 0.7692 0.8633 0.9778 0.9962 0.9869 0.7176 0.8243 0.7673
ConvNeXt[48] 196.24 0.8954 0.9067 0.8954 0.8976 0.7598 0.7452 0.7524 0.5985 0.6172 0.6077 0.9348 0.8884 0.9110 0.9877 0.9877 0.9877 1.0000 0.9900 0.9950 0.9759 0.9798 0.9779 0.9890 0.7692 0.8654 0.9813 0.9887 0.9850 0.6321 0.9054 0.7444
@@ 4 Experiments
We tested our model on a composite 9-class dataset. We created this dataset from three public sources: a brain tumor MRI dataset from Kaggle [49], a breast cancer X-ray dataset from Kaggle [50], and a chest X-ray pneumonia dataset from Mendeley Data [51] proposed by Daniel S. Kermany et al. [52]. This diverse dataset tests the model’s ability to work with different modalities and organs. This is an important capability for a general-purpose triage system. Table 1 shows the data distribution.
For a fair and rigorous comparison, all models were trained and evaluated under identical conditions. The baseline models were fine-tuned from their official ImageNet pretrained weights. The CNN backbones in our own model (DenseNet-121) were also initialized with ImageNet weights. This ensures that we are comparing the architectural advantages, not the benefits of pretraining.
4.1 Comparison with State-of-the-Art Baselines
A key part of our validation is comparing EVM-Fusion to strong, modern models. We chose a diverse set of baselines. Each one represents a different major architectural philosophy in computer vision. Table 3 shows the results of this comparison, including overall and per-class metrics.
subsectionAblation Studies: Validating the Synergistic Framework Our core hypothesis is that the strength of EVM-Fusion comes from the synergy between its components. To prove this, we conducted detailed ablation studies. The results are in Table 2. These studies were designed to answer two critical questions:
Is each feature extraction path necessary?: We tested the full model against versions where one or more of the paths were removed (e.g., EVM-Fusion (DHF) which removes the U-Net path).
Is our two-stage fusion mechanism effective?: We compared our full fusion module to simpler alternatives. These include Simple-Concat, Simple-Mean, CMA-Only, and NAF-Only.
@@ 5 Discussion
The results in Table 3 show that EVM-Fusion achieves state-of-the-art performance with an accuracy of 94.79%. Our model consistently outperforms all selected baseline models. These baselines represent the best of different architectural families. This result suggests that for the complex task of holistic image triage, a single design philosophy is not enough. The success of EVM-Fusion comes from its ability to effectively integrate these different approaches into a single, synergistic framework.
The ablation studies in Table 2 provide direct evidence for this synergy hypothesis. First, removing any of the three feature paths leads to a significant drop in performance. For example, removing the U-Net path (EVM-Fusion (DHF)) causes the accuracy to fall from 94.79% to 88.20%. This shows that the multi-scale features from the U-Net path provide unique and essential information.
Second, the fusion mechanism ablation is also very telling. The Simple-Concat baseline performs poorly (87.21% accuracy). This confirms that just combining the features is not an effective strategy. Using CMA-Only or NAF-Only improves the result, but the full two-stage system (EVM-Fusion (DUHF)) is clearly the best. This validates our central claim. The best way to fuse these heterogeneous features is to first use CMA for contextualization, then use the powerful NAF block to learn a complex fusion algorithm.
@@
Figure 2: Explainable AI (XAI) Analysis of Representative Correctly Classified Test Samples by EVM-DUHF Across All Classes. This figure showcases the intrinsic explainability outputs for nine correctly classified samples, one from each of the distinct medical image classes in the test set. Noted from (a) to (i). Similar to Fig 3, each sample is depicted through eight XAI visualizations. The analysis of these diverse, correctly classified instances aims to demonstrate the model’s adaptive behavior, the varying contributions of its different pathways, and its ability to focus on salient, class-relevant features across multiple organ types and imaging modalities.
Our XAI visualizations in Fig 2 provide a qualitative look at how this synergy works. Vision Mamba’s Adaptive Processing ($\Delta$-value Maps): The $\Delta$-value maps from the Vision Mamba modules reveal dynamic, input-dependent information gating. Higher $\Delta$ values (warmer colors) signify a greater emphasis on current patch information, indicating regions of novelty or high salience. Conversely, lower $\Delta$ values (cooler colors) suggest reliance on previously propagated contextual states, often where information is redundant or predictable. For instance, in a correctly classified ”normal” chest X-ray (Fig 2g), the DenseNet path’s Vim $\Delta$ maps show adaptive focus: the forward scan emphasizes lung boundaries upon encountering them, while the backward scan might highlight different boundary segments. The U-Net path’s Vim modules on the same sample might show more detailed, localized $\Delta$ peaks within the lung fields. Comparing forward and backward $\Delta$ maps (e.g., Fig 2b, a malignant tumor) often reveals differing ”novelty” assessments based on the directional context of Mamba’s selective scan, showcasing its capacity for nuanced contextual understanding. Path-Specific Spatial Attention: The spatial attention maps for the DenseNet and U-Net Mamba paths highlight prioritized image regions before global feature aggregation. In a correctly identified malignant tumor (Fig 2b), these maps focus on the breast tissue, with DenseNet potentially emphasizing broader areas while U-Net targets more specific regions like the nipple, possibly indicating sensitivity to microcalcifications or lesion-dense areas. Interestingly, Fig 2b also reveals a potential U-Net flaw, with some attention on blank image edges, underscoring the value of multi-path fusion to mitigate individual path imperfections. Traditional Feature Importance (SE-Attention): The Squeeze-and-Excitation (SE) attention weights within the traditional feature path quantify the contribution of specific GLCM and LBP features. For a correctly classified ”normal” lung (Fig 2g), features indicative of textural uniformity (LBP_bin_5 and LBP_bin_6) are often highly weighted. This demonstrates the model’s ability to leverage well-understood handcrafted features that complement deep learned representations, capturing intrinsic textural properties relevant to the diagnostic class. Cross-Modal Attention: A cornerstone of EVM-Fusion’s adaptability is the dynamic nature of its cross-modal attention weights, which vary significantly across different input images. For the benign tumors (Fig 2a), all paths show increased attention towards the U-Net path, suggesting a reliance on its detailed local information for tumor characterization. Conversely, for a ’healthy’ brain MRI (Fig 2e), the attention shifts, with U-Net drawing more from the traditional path if specific textures are key. This input-dependent, flexible weighting allows EVM-Fusion to adaptively modulate the influence of each feature pathway prior to the final Neural Algorithmic Fusion (NAF) stage. Notably, across many samples in Fig 2, most paths often receive substantial attention, indicating their collective importance and complementary roles. This inherent dynamism and balanced contribution underpin the superior performance of the full 3-path EVM-Fusion (EVM-DUHF) over ablated versions lacking one or more paths, as it allows the model to maximize information gain from diverse feature types and reduce over-reliance on any single potentially fallible source.
Figure 3: Explainable AI (XAI) Analysis of Representative Misclassified Test Samples by EVM-Fusion. This figure presents a consolidated view of intrinsic explainability outputs for five misclassified samples from the 9-class test set. Noted from (a) to (e). Each sample is analyzed using eight distinct XAI visualizations. From top to bottom, from left to right, they are DenseNet Vision Mamba Forward $\Delta$ Map, DenseNet Vision Mamba Backward $\Delta$ Map, U-Net Vision Mamba Forward $\Delta$ Map, U-Net Vision Mamba Backward $\Delta$ Map, DenseNet Path Spatial Attention, U-Net Path Spatial Attention, Traditional Feature Importance (SE Attention bar plot), Cross-Modal Attention Weights (heatmap). These visualizations aim to elucidate the model’s internal feature prioritization and fusion dynamics in instances of incorrect classification, providing insights into potential failure modes and areas of confusion.
@@
For (a) in Fig 3 both DenseNet and U-Net only focus on the same area of the breast. The model judges the malignant tumor (1) as benign (0), which indicates that there may be no tumor in this area. For (b) and (d) in Fig 3, it can be found that the DenseNet path focuses on the lower right corner, and U-Net pays almost no attention to the human body. This causes the model to even make cross-category errors, mistaking photos of the brain and chest for breasts. And (c) and (e) in Fig 3 also make the same mistake. Neither DenseNet nor U-Net pays good attention to the key organ - the lungs.